OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find 2·f(4). please help!!

OpenStudy (anonymous):

2.f(4)=2.(4^2+1)=2.(16+1)=2.17=34

OpenStudy (anonymous):

oh my god thank you so much!